If the circle 2x^2+2y^2=5x touches the line 3x+4y=k,find the values ok k.

Grade:11

1 Answers

Arun
5 years ago
Dear student
 
x² + y² - 5/4 x = 0     centre = (5/8, 0)
Line : 3x + 4y - k = 0
 
If this tangent then perpendicular length from centr centre be equal to radius
 
|15/8 - k| /5 = 5/8
 
Now opem this and you will get two vues of k